Given one equation about pair of lines like aX2+Y2+4X+3Y=0 FIND VALUE OF a
 

Grade:12th pass

1 Answers

Vikas TU
5 years ago
Dear Student,
The general equation of a pair of straight lines is ax2+2hxy+by2+2gx+2fy+c=0.
The condition that this represents a pair of straight lines is (hy – af)2 = (h2 – ab)(g2 – ac).
In this equation, we have  ax2+y2+4x+3y=0.
Therefore we have h=0, b=1, g=2, f=3/2, c=0.
Substituting this in the equation, we get (- 3a/2)2 = (0 – a)(4 – 0) = - 4a.
Therefore 9a2/4 = - 4a. 
a = 0 or a = - 16/9.
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)